OpenStudy (anonymous):

A blimp flying at an altitude of 500 feet lies directly over a line from soldier field to the adler planetarium on Lake Michigan. The angle of depression from the blimp to the stadium is 32 degrees and from the blimp to the planetarium is 23 degrees. Find the distance the blimp would have to fly in order to land in Soldier field.

OpenStudy (anonymous):

 tan 23=500/a a=500/tan 23 = 1177.93' tan 32=500/b b=500/tan 32 = 312.43' 1177.93'+312.43'=1490.36'